One number is six less than another. Their sum is twelve. Find the numbers. (Enter solutions from smallest to largest.)
Answer by ewatrrr(24785)


Hi
One number is six less than another
Let x and (x-6) represent the two numbers
Question states***
x + (x-6) = 12
Solving for x
2x = 18
x = 9
(x-6)= 3 The two numbers are 3 and 9.
CHECKING our Answer***
3 + 9 = 12